Question 14 4 points Save Answer You are pushing a 3.15-kg box toward the right as it moves, in the direction of your push, over a patch of rough grass. If the magnitude of your push is 18 4 N, and the coefficient of kinetic friction with the grass is 0.350, what will be the magnitude and direction of the acceleration of the box? O a.-2.4 mis2 b. -1.2 m/s2 O c. 1.2 m/s2 d.2.4 m/s2
Answer #1

Solution)

By Newton's law

We know, F=ma

A=F/m

Fnet=F-u*mg

So, a=F-umg/m

A=18.4 N-0.35*3.15*9.8/3.15 =2.41 m/s^2

=====

Good luck!:)